OpenStudy (anonymous):

Find the derivative of the funtion. f(x)=x^2(x-2)^4

OpenStudy (anonymous):

\[f(x)=x ^{2}(x-2)^{4}\]

sam (.sam.):

Use product rule, y' = (x^2)(4)(x-2)^3+(x-2)^4 (2x) \[4x^{2}(x-2)^{3}+2x(x-2)^{4}\]

sam (.sam.):

Teach you one technique, y=(left)(right) y'=(copy left)(differentiate right)+(copy right)(differentiate left)

OpenStudy (anonymous):

yea product rule

OpenStudy (anonymous):

you get 2x(x-2)^4+x^2(4(x-2)^3)

OpenStudy (anonymous):

then apply the chain rule which i do not understand

sam (.sam.):

Noneed to use chain rule here.

OpenStudy (anonymous):

the answer in the back of the book says

OpenStudy (anonymous):

\[x ^{2}[4(x-2)^{3}(1)] + (x-2)^{4}(2x) = 2x(x-2)^{3}(3x-2)\]

OpenStudy (anonymous):

i dont understand hgow to get to 2x(x-2)^3(3x-2)

OpenStudy (anonymous):

f(x)=x^2(x-2)^4 split these things up durp, g(x) = x^(2) g'(x) = 2x d(x) = (x-2)^4 USE CHAIN RULE!!! s(x) = x^4 s'(x) = 4x^(3) j(x) = x-2 j'(x) = 1 so d'(x) =4(x-2)^(3)*1 so Now product rule g(x) = x^(2) g'(x) = 2x d(x) = (x-2)^4 d'(x) =4(x-2)^(3)*1 2x((x-2)^4) + (4(x-2)^(3))x^(2)

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

f'(x) = 2x((x-2)^4) + (4(x-2)^(3))x^(2) Best way to deal with derivatives when you are first learning them is to split them into different functions and derive them then put them all together

OpenStudy (anonymous):

After you do this about a 1,000 times you will be able to do it in your head

OpenStudy (anonymous):

Any questions? also if you do it in the order I do it you wont mess up using Quotient rule

OpenStudy (anonymous):

sorry let me read what you wrote

OpenStudy (anonymous):

had stepped away

OpenStudy (anonymous):

Remember Chain rule is just g(x) = x^4 g'(x) = 4x^(3) s(x) = x-2 s'(x) = 1 we throw away g(x) and sub s(x) into g'(x) and multiply by s(x) derivative so g'(s(x)) * s'(x)

OpenStudy (anonymous):

this is the fool proof way of doing derivatives when you first start out at least it was for me

OpenStudy (anonymous):

ok let me try to reword this

OpenStudy (anonymous):

x^2(x+2)^4

OpenStudy (anonymous):

f(x)=x^2(x-2)^4 f'(x) = 2x *(x-2)^4 + 4x^2 (x -2) ^3 = (x - 3) ^3 [ 2x(x-2) + 4x^2 ] = (x - 3) ^3 [ 6x^2 - 4x ]

OpenStudy (anonymous):

(2x)((x+2)^4) + 4x^2(x-2)^3

OpenStudy (anonymous):

g(x)d(x) Take derivative of g(x) take derivative of d(x) you have to use chain rule to do so g'(x)d(x) + g(x)d'(x) if you do it in this order it works for quient rule as well g(x)/d(x) (g'(x)(d(x)) - g(x)d'(x))/(d(x))^(2)

OpenStudy (anonymous):

ah so i then take (2x)((x+2)^4) + 4x^2(x-2)^3 and split it like [2x(x+2)^4 +4x^2](x-3)^3 ?

OpenStudy (anonymous):

and do the product rule again?

OpenStudy (anonymous):

no!

OpenStudy (anonymous):

f'(x) = 2x((x-2)^4) + (4(x-2)^(3))x^(2) That is it you can simplify it but that is the derivative

OpenStudy (anonymous):

if you wanted to take the second derivative you would use product rule again.

OpenStudy (anonymous):

ok so i need to simplify f'(x) = 2x((x-2)^4) + (4(x-2)^(3))x^(2)

OpenStudy (anonymous):

its only asking for the first derivative so yeah. Is there anything you do not understand about the method I showed you? If you have to you can

OpenStudy (anonymous):

so i factor out (x-2)^3

OpenStudy (anonymous):

you can also factor out 2x

OpenStudy (anonymous):

2x(x-2)^(3)(x - 2 + 2x)

OpenStudy (anonymous):

so i would end up with 2x(x-2)^3(x-2+2x)

OpenStudy (anonymous):

2x(3x-2)(x-2)^(3)

OpenStudy (anonymous):

2x^(1) and x^(1) can be added together

OpenStudy (anonymous):

you have been so much help

OpenStudy (anonymous):

i had the right answer the whole time just didnt see the factor

OpenStudy (anonymous):

because they are to the same power, remember they have to be to the same power to add them you can't add x^(2) + x^(1) or x^(3) + x Not to be partrionizing

OpenStudy (anonymous):

*patronizing

OpenStudy (anonymous):

yeah math can be a hassel sometimes lol

OpenStudy (anonymous):

no i understand now

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

np take it easy